OpenStudy (anonymous):

Using the power, product, and quotient rules to simplify expressions with negative exponents: (u^-2/ w^3)(-2u^2v^3w)^-3

OpenStudy (anonymous):

\[\frac{u^{-2}}{w^3}(-2u^2v^3w)^{-3}\]

OpenStudy (anonymous):

Someone going to do this?

OpenStudy (anonymous):

How Did You Write It Like That ?

OpenStudy (anonymous):

\frac{u^{-2}}{w^3}(-2u^2v^3w)^{-3}

OpenStudy (anonymous):

You put that in the equation box..

OpenStudy (anonymous):

you are starting with \[\frac{u^2}{w^3}\left(-2u^2v^3w\right)^3\] which if we deal with the large bracket cube, this then is equal to\[\frac{u^2}{w^3}\left((-2)^3(u^2)^3 (v^3)^3(w)^3\right)\] or put more simply \[\frac{u^2}{w^3}\left(-8u^6v^9w^3\right)\] as (-2)x(-2)x(-2)=-8, and \[(u^2)^3=u^2*u^2*u^2=u^{(2+2+2)}=u^6\] and so on.Now you also have \[\frac{1}{w^3}=w^{-3}\] so your equation is now \[u^2w^{-3}\left(-8u^6v^9w^3\right)=-8u^{2+6}v^9w^{3-3}=-8u^8v^9\]